# Optimizing Inventory Placement for a Downstream Online Matching Problem

### v1

@@ 1.1 Theoretical Results and Techniques
Our guarantee is strictly better than the $(1-1/e)/2$-approximation implied by Bai et al. (2022b) for Fluid Placement (their guarantee is stated as 1/4 and holds in a generalized setting; it improves to $(1-1/e)/2$ in the setting stated here). Our result also provides theoretical justification for why Offline appears to outperform Fluid Placement in the experiments of DeValve et al. (2023b). As $d\to\infty$, our guarantee decreases to approach their guarantee of $(1-1/e)/2$, which is obtained by using submodular optimization to approximately solve Fluid Placement within a factor of $1-1/e$. However, the submodular optimization approach cannot improve (see Section 7) as $d$ gets small, which is often the case in practice—$d=2$ in the celebrated long-chain network (Jordan and Graves 1995) or the JD.com network (Shen et al. 2020). This is why we take a randomized rounding approach instead, and moreover use it to analyze Offline Placement, since we cannot use integrality properties about demand in Fluid Placement. Another benefit of using Offline Fulfillment is that in some arrival models of online matching, the fluid relaxation is too loose to yield constant-factor competitive ratios for online algorithms, as is the case in the Spatial Independence model (Aouad and Ma 2022, Proposition 1).
@@ 1.2 Experimental Results on Public Data
Information Required
Medium (all moments of aggregate demand for each type)
High (exact demand arrival sequence)
Low (first moment of aggregate demand for each type)
Computation Time
Requires sampling
Requires simulation
Fastest (still no poly-time algorithm)
Surrogate Approximation
$1-(1-1/d)^{d}$ $\boldsymbol{(*)}$
(LP and rounding)
$1-1/e$ (Bai et al. 2022b)
(Submodular optimization)
Overall Approximation
$(1-(1-1/d)^{d})/2$ $\boldsymbol{(*)}$
(Under both demand models)
$(1-1/e)/2$ (Bai et al. 2022b)
(Only under Temporal Independence)
Empirical Justification
Beats $\mathsf{FLU}$ (DeValve et al. 2023b)
Beats $\mathsf{MYO}$ given a good fulfillment policy $\boldsymbol{(*)}$
Beats $\mathsf{FLU}$ $\boldsymbol{(*)}$
@@ 4.2 Approximating the Surrogate: Randomized Rounding
One possible approach for approximating $\max_{x_{\in}\mathcal{X}}\mathsf{OFF}(x)$ is to use submodular optimization techniques. However, this approach fails to get an approximation that improves for small $d$. Bai et al. (2022b) show that $\mathsf{OFF}(x,D)$ is a submodular function over $x\in\mathcal{X}$ for any fixed $D$. Since submodularity is preserved under a convex combination, $\mathsf{OFF}(x)$ is also a submodular function over $x\in\mathcal{X}$. It is straightforward to verify that $\mathsf{OFF}(x)$ is also a non-decreasing function of $x$, implying that we could obtain a $(1-1/e)$-approximation of $\max_{x\in\mathcal{X}}\mathsf{OFF}(x)$ by constructing an inventory placement greedily (Calinescu et al. 2007). More precisely, the submodular approach would yield an approximation of $(1-(1-1/Q)^{Q})$, where $Q$ is the number of inventory units to be placed. In applications $Q$ is large, so this approximation would be close to $1-1/e$. This factor cannot be further improved even if $d$ is small, as shown in Section 7.
@@
The proof of this lemma is deferred to Subsection 6.2. Equipped with Lemma 4.5, we now lower-bound the reward that can be obtained from each demand type, in expectation over the randomized placement and assignment.
@@
The proof of this lemma is deferred to Subsection 6.3. With these results, we now complete the proof of Theorem 4.3, presented in Subsection 6.4.
@@
The proof of this lemma is deferred to Subsection 6.5.
@@ 4.3 Approximating the Surrogate: Sample-average Approximation
where $(\sigma_{k})_{k\in[K]}$ are independent Rademacher random variables, defined as random variables that are either $+1$ or $-1$, with probability $1/2$ each. We note that in (10) we are normalizing all values $\mathsf{OFF}(x,D^{k})$ to lie in [0,1] (all rewards $r_{ij}$ lie in [0,1] and we have only $Q$ units, so the total reward is upper-bounded by $Q$) and subtracting it from 1 to get a loss function, to obtain the following result, whose proof is deferred to Subsection 6.6.
@@
The proof of this lemma is deferred to Subsection 6.7. With this we are finally able to write the proof of Theorem 4.8, which is presented in Subsection 6.8.
@@ 4.4 Approximation Algorithm and Main Result
The proof of this theorem is deferred to Subsection 6.9.

### v2

@@ 1.1 Theoretical Results and Techniques
Our guarantee is strictly better than the $(1-1/e)/2$-approximation implied by Bai et al. (2022b) for Fluid Placement (their guarantee is stated as 1/4 and holds in a generalized setting; it improves to $(1-1/e)/2$ in the setting stated here). Our result also provides theoretical justification for why Offline appears to outperform Fluid Placement in the experiments of DeValve et al. (2023b). As $d\to\infty$, our guarantee decreases to approach their guarantee of $(1-1/e)/2$, which is obtained by using submodular optimization to approximately solve Fluid Placement within a factor of $1-1/e$. However, the submodular optimization approach cannot improve (see LABEL:sec:greedySucks) as $d$ gets small, which is often the case in practice—$d=2$ in the celebrated long-chain network (Jordan and Graves 1995) or the JD.com network (Shen et al. 2020). This is why we take a randomized rounding approach instead, and moreover use it to analyze Offline Placement, since we cannot use integrality properties about demand in Fluid Placement. Another benefit of using Offline Fulfillment is that in some arrival models of online matching, the fluid relaxation is too loose to yield constant-factor competitive ratios for online algorithms, as is the case in the Spatial Independence model (Aouad and Ma 2022, Proposition 1).
@@ 1.2 Experimental Results on Public Data
Information Required Medium (all moments of aggregate demand for each type) High (exact demand arrival sequence) Low (first moment of aggregate demand for each type)
Computation Time Requires sampling Requires simulation Fastest (still no poly-time algorithm)
Surrogate Approximation $1-(1-1/d)^{d}$ $\boldsymbol{(*)}$ (LP and rounding) $1-1/e$ (Bai et al. 2022b) (Submodular optimization)
Overall Approximation $(1-(1-1/d)^{d})/2$ $\boldsymbol{(*)}$ (Under both demand models) $(1-1/e)/2$ (Bai et al. 2022b) (Only under Temporal Independence)
Empirical Justification Beats $\mathsf{FLU}$ (DeValve et al. 2023b) Beats $\mathsf{MYO}$ given a good fulfillment policy $\boldsymbol{(*)}$ Beats $\mathsf{FLU}$ $\boldsymbol{(*)}$
@@ 4.2 Approximating the Surrogate: Randomized Rounding
One possible approach for approximating $\max_{x_{\in}\mathcal{X}}\mathsf{OFF}(x)$ is to use submodular optimization techniques. However, this approach fails to get an approximation that improves for small $d$. Bai et al. (2022b) show that $\mathsf{OFF}(x,D)$ is a submodular function over $x\in\mathcal{X}$ for any fixed $D$. Since submodularity is preserved under a convex combination, $\mathsf{OFF}(x)$ is also a submodular function over $x\in\mathcal{X}$. It is straightforward to verify that $\mathsf{OFF}(x)$ is also a non-decreasing function of $x$, implying that we could obtain a $(1-1/e)$-approximation of $\max_{x\in\mathcal{X}}\mathsf{OFF}(x)$ by constructing an inventory placement greedily (Calinescu et al. 2007). More precisely, the submodular approach would yield an approximation of $(1-(1-1/Q)^{Q})$, where $Q$ is the number of inventory units to be placed. In applications $Q$ is large, so this approximation would be close to $1-1/e$. This factor cannot be further improved even if $d$ is small, as shown in LABEL:sec:greedySucks.
@@
The proof of this lemma is deferred to LABEL:app:proof_neg_corr. Equipped with Lemma 4.5, we now lower-bound the reward that can be obtained from each demand type, in expectation over the randomized placement and assignment.
@@
In LABEL:app:proof_Zbound we prove Lemma 4.6, noting that a similar proof has appeared in Brubach et al. (2021). With these results, we now complete the proof of Theorem 4.3, presented in LABEL:app:proof_rounding.
@@
The proof of this lemma is deferred to LABEL:app:tight_example.
@@ 4.3 Approximating the Surrogate: Sample-average Approximation
where $(\sigma_{k})_{k\in[K]}$ are independent Rademacher random variables, defined as random variables that are either $+1$ or $-1$, with probability $1/2$ each. We note that in (10) we are normalizing all values $\mathsf{OFF}(x,D^{k})$ to lie in [0,1] (all rewards $r_{ij}$ lie in [0,1] and we have only $Q$ units, so the total reward is upper-bounded by $Q$) and subtracting it from 1 to get a loss function, to obtain the following result, whose proof is deferred to LABEL:app:proof_rad_lem.
@@
The proof of this lemma is deferred to LABEL:app:proof_lipschitz. With this we are finally able to write the proof of Theorem 4.8, which is presented in LABEL:app:proof_stat_learn_thm.
@@ 4.4 Approximation Algorithm and Main Result
The proof of this theorem is deferred to LABEL:app:proof_main.
@@ 5.6 Results
[…]
The authors sincerely thank anonymous reviewers for identifying an error in the original proof of Lemma 4.4. The statement is correct and the proof has since been fixed.